OpenStudy (anonymous):

Solve using the substitution method. Show your work. If the system has no solution or an infinite number of solutions, state this. 4x + 4y = -36 4x + 4y = 12

Parth (parthkohli):

4x = -36 - 4y So, (-36 - 4y) + 4y = 12 -36 = 12 This system has no solution

OpenStudy (anonymous):

x + y = -9 (x + y = 3) *-1 x + y = -9 -x - y = -3 =-12 so it has no solution :D

OpenStudy (anonymous):

Thank you very much!